APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday June 4, 2004 (MJD 53161) ---=== OBSERVING TEAM ===--- Atsuko Nitta ---=== OBSERVING PLAN ===--- Pre-shake checkout. ---=== OBSERVING SUMMARY ===--- Starting the pre-shake down work. o Cleaned out /data/photo, astro, spectro. o Copied the /astrolog/BRIGHT_TIME_MJDs to /skycat. o Cleaned tape drives. o New fiducial table v1_20 built and checked in. MCP rebooted to make sure it reads the new tables. Tomorrow, I will do the spectro monthly checkout. ---=== OBSERVING LOG ===--- Copied the /astrolog/53151 to 53160 to /skycat/oplogs/astrolog. Then deleted /astrolog/53151to 53159. Cleaned out /data/phot, /data/astro. Claned out /data/spectro on host and sos except for 53155 directory which contains data contained by Connie. Fiducial tables built and checked in. New fiducial is v1_20. MCP was rebooted ensure that it read in the new table. Cleaned the DA tape drives. JB told me that he changed the PT FF lamp this afternoon and took flat field frames to ensure it is working well. Steve found right before he left taht the pressure on the imager dewar was less than 20psi. So he swapped the dewar to a one with much higher pressure (~30psi) before going home. I will do the spectroMonthly checkout tomorrow. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 27009, plate -9999 ------- 01:58 27009 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== DATA TAPE SUMMARY ===--- Goes: JL Stays: JL ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:46Z 64 31 5 177 (S) 1449 - 23:17Z 66 34 1 193 (SSW) 2239 - 23:50Z 65 36 6 267 (W) 3200 - 00:22Z 64 37 9 316 (NW) 2850 - 00:54Z 64 39 13 243 (WSW) 3287 - 01:26Z 65 36 6 264 (W) 2322 - 01:59Z 62 35 8 232 (SW) 2187 - 02:29Z 60 33 8 272 (W) 3518 - 03:02Z 59 31 11 321 (NW) 3237 - ---=== TELESCOPE STATUS ===--- ---=== SOFTWARE USED ===--- IOP/SOP: v3_138_0 -- test version from shakeup Watcher: v2_29_0 MCP: v5_25_0 New fiducial table v1_20 TPM: tpm_v2_42_0 AstroDa: v14_47 TCC: TCC 2.7.2 November 11 2003 sdssProcedures: v1_90 SoS: v4_10_7 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- ---=== PROBLEMS IN DETAIL ===--- Today's TPM error in watcher: dewar (Spectro Dewar Scale Communications Fault: 2004-06-05 02:50:50Z)